OpenStudy (anonymous):

A boy weighing 45 Kg stands on a spring balance that is going down with a constant speed of 3m/s if the lift is brought to halt by a constant deceleration, in a distance of 10m, the scale reading during this period is?

OpenStudy (anonymous):

@shubhamsrg @ajprincess @hartnn @ujjwal

OpenStudy (anonymous):

V^2 - u^2 = 2as -9 = 2*a*10 a = -9/10

OpenStudy (anonymous):

@mukushla

OpenStudy (anonymous):

R1 = m(g-a)

OpenStudy (anonymous):

sorry im busy

OpenStudy (anonymous):

man..just a minute plzz

OpenStudy (ajprincess):